OpenStudy (anonymous):

I have troubles with this question. Would you mind answering it? And make me understand this. Just answer TRUE or FALSE 1.) A growth rate of 20% corresponds to a growth factor of 1.02. 2.) A decay of 0.675 corresponds to a decay rate of 3.25%. 3.) Quantities that increase by a constant amount represents linear growth, whereas quantities that increase by a constant percent represents exponential growth.

OpenStudy (skullpatrol):

1) false

OpenStudy (skullpatrol):

questions?

OpenStudy (anonymous):

i mean statements. Sorry I'm very stress now

OpenStudy (skullpatrol):

do you have any questions about the answer to #1?

OpenStudy (anonymous):

Yes. How did you know that it was false?

OpenStudy (skullpatrol):

20% = 20/100 = 0.20

OpenStudy (skullpatrol):

$$\Huge % =\frac{1}{100}$$

OpenStudy (skullpatrol):

Is latex broken?

OpenStudy (skullpatrol):